# Investigation of hydrogen under-expanded jets in gaseous propulsion systems

### v1

@@ 5.1 Under-expanded H2 and Air jets
We can explain the differences between the two jets by examining the vorticity equation
@@
where we have the material of the vorticity on the left-hand side of 7, stretching and tilting in the first two terms on the right-hand side, the baroclinic term (the third term on the right), and the diffusion by viscosity in the last term. The baroclinic term produces vorticity when the density gradient is not aligned with the pressure gradient; this term is always present in compressible flows, but we expect strong effects when injecting light gas (hydrogen, in the present simulations) into heavy environment gas with respect to the situation of air-in-air or methane-in-air jets (Duronio & De Vita, 2024). Figure 9 reports the component of this term in non-dimensional form for the validation case 2 and R1 jets.
@@
For the H2-in-air jet, the pressure gradient is almost parallel to the axis and directed to the left, while the density gradient is pointing outward with respect to the axis because its core is occupied by hydrogen, as shown by the arrows on the right of figure 9; therefore, the baroclinic term in the vorticity equation induces counterclockwise rotation in the upper part of the figure and clockwise rotation in the lower portion (as shown by the circular arrows in the picture) that push the hydrogen outward. On the contrary, in the air-in-air flow, the density gradient is pointing toward the jet axis, and consequently, the induced rotation is opposite to the one observed with the hydrogen flow. In addition, the magnitude of this term is much lower.
This baroclinic term justifies the completely different morphology of the outer jet between the single-species and the two-species flow. Nonetheless, the shock structure in the core is remarkably similar. Figure 10 is a zoomed view of the Mach disc for the two investigated cases performed after .
@@
The figure shows that the structure of the primary shock cell is not affected by the species injected, as both the converging shocks and the normal shock have a similar morphology. As previously reported, the two-species jet exhibits a broader shear layer outside the jet core, where the hydrogen quickly mixes with the surrounding air. To quantify this aspect, we reported in figure 11 the magnitude of the vorticity vector after . The plot is a volumetric rendering where we removed a quarter of the jet to show its internal structure.
The vorticity is clearly higher for the hydrogen jet. In the near nozzle zone, both the jets present streamwise vortices outside the potential core with hairpin vortices downstream. As already underlined, these aspects are relevant in chemically reactive flows since the efficiency of the subsequent combustion process strongly depends on the mixing characteristics.
5.2 Effects of the pressure ratio on the H2 jets
@@
The early stages of the jets significantly differ when increasing the pressure ratio. The jet with the lower pressure ratio shows two converging shocks that merge at a point located on the jet’s axis and a pronounced primary vortex ring. When increasing the upstream pressure, the classical Mach disc shock appears, and the structure of the vortex rings changes. We can observe a primary vortex ring and a significant secondary vortex ring. A cap flow discontinuity delimits a distinguishable zone downstream of the Mach disc shock. The pressure ratio also has a significant effect on the shear layer. Indeed, when looking at the snapshots in figure 12, it can be concluded that a higher injection pressure promotes a broader shear layer, enhances the Coanda effect, and, as discussed below, helps form an ignitable air/fuel mixture. At t = (figure 13), the final under-expanded structures that characterized the jet can be observed.
Following the classification proposed by Franquet et al. (2015) and Duronio et al. (2023), the jet of case R1 is highly under-expanded, while the jets R2 and R3 are extremely (or very highly) under-expanded. Indeed, in the first case, the jet has a “barrel” structure that repeats several times downstream. The triple point can be recognized as the intersection of the intercepting shock, the Mach disc, and the reflected shock. Increasing the injection pressure, the structure of the jet is dominated by a single barrel shock. The Mach disc is no longer characterized by a normal shock, but a relevant curvature appears. The lack of a normal shock sequence promotes jet penetration and reduces the so-called potential core’s length, increasing the mixing between hydrogen and air; this aspect is shown in figure 14, where it can be observed that the zone of the jet characterized by the sequence of barrel shocks does not show any mixing activity, which is completely inhibited. The increase in the pressure ratio reduces the length of this zone.
@@
Obtaining a quantitative evaluation of the overall quality of the mixture is essential when dealing with combustion processes. Thus, we defined a discrete mass-weighted probability density function () as:
@@
The PDFs for the three cases are very similar, showing a peak around ; of course, the amount of hydrogen mass injected grows with the injection pressure.
5.3 Effects of the nozzle shape
The nozzle design for gaseous injectors is essential to developing efficient injection devices. For this reason, we investigate two annular nozzles characterized by jet angles equal to 90∘ (case A4) and 135∘ (case A5). Figure 17 compares the jet’s transient evolution for the early stages of the simulated transient. As before, we plotted the .
The evolution is very different when compared with the round jet. At t =, the hollow-cone jets present a conical structure characterized by the angle imposed. The pressure inside the cone becomes gradually lower than the external one, and the jet continues its evolution by contracting towards the x-axis, as shown in figure 18 where we reported the time-averaged velocity vectors and the pressure field over on the axial plane.
Outside the core, in the near nozzle zone, air cannot enter within the jet due to the presence of a series of shock cells. On the other hand, we can observe a series of vortices rotating toward the central axis that drags the hydrogen into the jet’s core. This flow path extends for approximately 4D downstream of the nozzle exit section. The intensity of this phenomenon increases with the jet cone angle. As illustrated by figure 19, this circulation remains active until the jet spreads in the surroundings.
The subsequent injection stages are characterized by the jet advancing toward the right boundary of the domain with a velocity greater for jet A4 because of the smaller cone angle. Although the shape of the jet is similar, the cone angle modifies the hydrogen concentration, as shown in the figure by . Various experimental investigations confirm this evolution mechanism, and it is characteristic of hollow cone jets Lee et al. (2021); Coratella et al. (2024); Montanaro, Alessandro et al. (2024).
@@ 5.3 Effects of the nozzle shape
The hollow conical structure of the jet determines a ring-shaped spatial arrangement of the shock waves. The single Mach disc observed with the circular nozzle disappears, but barrel shocks can also be recognized for test cases in A4 and A5; moreover, their size diminishes with the increasing jet angle. The number of these shock cells is larger for case A5. We can conclude that, when performing a hollow cone injection, the increase of the jet angle reduces the Mach number and the hydrogen velocity; therefore, with the same nozzle area, the injection of the same amount of requires a longer time.
To quantitatively assess the mixture quality obtainable with a hollow cone injection, in figure 21 we plotted the PDF and the average concentration into the domain as a function of time for the cases R2, A4, and A5. To have comparable results, we plotted the PDF at the time instant required to have an injected non-dimensional mass equal to 100 for all cases.
@@
The increased jet angle promotes the mixing with the surrounding air and almost completely removes the hydrogen-rich potential core. It can be seen that the region where the hydrogen mass fraction is close to one is shrunk, for case A5, to a hollow cone around the annular inflow section.
6 Conclusions
We investigated hydrogen under-expanded jets using Large Eddy Simulations of round and annular nozzles. We focused on the jet’s characteristics and, more importantly, the mixture formation process.
The simulations were verified and validated against experimental particle image velocimetry and Schlieren images when available. The simulated jets, the Mach disc dimensions, and the resulting velocity field closely align with experimental observations.
The analysis covers injection pressure ratios from 4.2 to 25.6 and examines the nozzle geometrical configuration’s effects on the air-fuel mixture’s characteristics. The main outcomes are:
The transient evolution of hydrogen under-expanded jets radically differs from the one observed for air. In particular, the mixing outside the jet core is enhanced by the baroclinic effect. On the contrary, the structure of the jet core is similar for the two flows in terms of shock structures.
The physical characteristics of the hydrogen trigger an intense interaction with the surrounding air. The hydrogen jet develops very high vorticity and engulfs a large air volume.
The pressure ratio is a key parameter in controlling the jet’s characteristics. It can be used to inject the proper amount of fuel in a reduced time.
Increasing the pressure ratio, the repeated shock cells disappear, and we have a unique Mach disc. Besides, the potential core becomes smaller for higher pressure ratios. As the potential core is the zone of the jet where there is no mixing activity when reducing it, the mixture quality improves.
Hollow cone jets have a drastically different morphology when compared to the classical round jets. Ring-shaped shock cells substitute the Mach disc, and the jet from highly under-expanded becomes weakly under-expanded. The Mach number decreases, and injecting the same amount of hydrogen requires a longer time.
The recirculation zone developing downstream the nozzle for approximately 4D significantly enhances the mixing activity, removes the hydrogen-rich core and gives a higher quality mixture.
A jet cone angle greater than 90∘ improves the mixture quality from the combustion point of view. With a jet cone angle equal to 135∘ we manage to quickly reduce the average hydrogen concentration already in the first phase of the injection and to remove the zones where the hydrogen concentration exceeds 0.4 completely.
Acknowledgment
This work has been funded by the European Union - NextGenerationEU, Mission 4, Component 1, under the Italian Ministry of University and Research (MUR) National Innovation Ecosystem grant ECS00000041 - VITALITY - CUP E13C22001060006.
Declaration of Interests
The authors report no conflict of interest.

### v2

@@ V.1 Under-expanded H2 and air jets
We can explain the differences between the two jets by examining the vorticity equation:
@@
where we have the material derivative of the vorticity on the left-hand side of equation 7, stretching and tilting in the first two terms on the right-hand side, the baroclinic term (the third term on the right), and the diffusion by viscosity in the last term. The baroclinic term produces vorticity when the density gradient is not aligned with the pressure gradient; this term is always present in compressible flows, but we expect strong effects when injecting light gas (hydrogen, in the present simulations) into a heavy gas environment, like air-in-air or methane-in-air jets Duronio and De Vita (2024). Figure 11 reports the component of this term in non-dimensional form for the validation case 2 and R1 jets.
@@
For the H2-in-air jet, the pressure gradient is almost parallel to the axis and directed to the left, while the density gradient is pointing outward because the jet core is occupied by (the lighter) hydrogen, as shown by the arrows on the right of figure 11; therefore, the baroclinic term in the vorticity equation induces counterclockwise rotation in the upper part of the figure and clockwise rotation in the lower portion (as shown by the circular arrows in the picture) that push the hydrogen outward. On the contrary, in the air-in-air flow, the density gradient points toward the axis; consequently, the induced rotation is opposite to the one observed with the hydrogen flow. In addition, the magnitude of this term is much lower. This baroclinic term justifies the completely different morphology of the outer jet between the single-species and the two-species flow. Nonetheless, the shock structure in the core is remarkably similar. Figure 12 is a zoomed view of the Mach disc for the two investigated cases performed after .
@@
The figure shows that the structure of the primary shock cell is not affected by the species injected, as both the converging shocks and the normal shock have a similar morphology. As previously reported, the two-species jet exhibits a broader shear layer outside the jet core, where the hydrogen quickly mixes with the surrounding air. To quantify this aspect, we reported in Figure 13 the magnitude of the vorticity vector after .
The plot is a volumetric rendering where we removed a quarter of the jet to show its internal structure. The vorticity is higher for the hydrogen jet. In the near nozzle zone, both the jets present streamwise vortices outside the potential core in the shear layer with small-scale vortices downstream. As already underlined, the vorticity field provides an overview of the turbulence which is relevant in chemically reactive flows since it drives the mixing of the air with the fuel and so the efficiency of the subsequent combustion process Buttay, Lehnasch, and Mura (2018); Hamzehloo and Aleiferis (2016b). A more detailed quantitative comparison of the two jets is reported in the figure 14, where the axial plots of non-dimensional density, pressure, and Mach number fields are reported.
As already discussed, relevant differences can be recognized in the density field.
V.2 Effects of the pressure ratio on the H2 jets
@@
The early stages of the jets significantly differ when increasing the pressure ratio. The jet with the lower pressure ratio shows two converging shocks that merge at a point on the jet’s axis and a pronounced primary vortex ring. When increasing the upstream pressure, the classical Mach disc shock appears, and the structure of the vortex rings changes. We can observe a primary vortex ring and a significant secondary vortex ring. A cap flow discontinuity delimits a distinguishable zone downstream of the Mach disc shock. The pressure ratio also has a significant effect on the shear layer. Indeed, looking at the snapshots in Figure 15, it can be concluded that a higher injection pressure promotes a broader shear layer, enhances the Coanda effect, and, as discussed below, helps form an ignitable air/fuel mixture. At t = (figure 16), the final under-expanded structures that characterized the jet can be observed.
Following the classification proposed by Franquet et al. (2015) and Duronio et al. Duronio, Villante, and De Vita (2023), the jet of case R1 is highly under-expanded, while the jets R2 and R3 are extremely (or very highly) under-expanded. Indeed, in the first case, the jet has a “barrel” structure that repeats three times downstream. The triple point can be recognised as the intersection of the intercepting shock, the Mach disc, and the reflected shock. When increasing the injection pressure, the structure of the jet is dominated by a single barrel shock. A normal shock no longer characterises the Mach disc, but a relevant curvature appears. The Mach disc’s width and height measurements are reported in Table 3 for the three pressure ratios investigated.
| Jet # | R1 | R2 | R3 |
| Height [-] | 1.17D | 2.25D | 3.3D |
| Width [-] | 0.2D | 1.06D | 1.5D |
The lack of a normal shock sequence promotes jet penetration and reduces the so-called potential core’s length, increasing the mixing between hydrogen and air; this aspect is shown in figure 17, where it can be observed that the zone of the jet characterized by the sequence of barrel shocks does not show any mixing activity, which is completely inhibited. The increase in the pressure ratio reduces the length of this zone.
@@
Obtaining a quantitative evaluation of the overall quality of the mixture is essential when dealing with combustion processes. Thus, we defined a discrete mass-weighted probability density function as:
@@
The PDFs for the three cases are very similar, showing a peak around ; the amount of hydrogen mass injected grows with the injection pressure.
V.3 Effects of the nozzle shape
The nozzle design for gaseous injectors is essential to developing efficient injection devices. For this reason, we investigate two annular nozzles characterised by jet angles equal to 90∘ (case A4) and 135∘ (case A5). Figure 20 compares the jet’s transient evolution for the early stages of the simulated transient. As before, we plotted the .
The evolution is very different when compared with the round jet. At t =, the hollow-cone jets present a conical structure characterized by the angle imposed. The pressure inside the cone becomes gradually lower than the external pressure. The jet continues its evolution by contracting towards the x-axis, as shown in figure 21, where we reported the time-averaged velocity vectors and the pressure field over on the axial plane.
Outside the core, in the near nozzle zone, air cannot enter within the jet due to a series of shock cells. On the other hand, we can observe a series of vortices rotating toward the central axis that drag the hydrogen into the jet’s core. This flow path extends for approximately 4D downstream of the nozzle exit section. The intensity of this phenomenon increases with the jet cone angle. Figure 22 reports a quantitative evaluation of the pressure diminishing.
The pressure reduction reaches 30% in the case A5, while in the case A4 it is at most 20%. As illustrated by Figure 23, these vortical structures remain active until the jet spreads in the surroundings.
The subsequent injection stages are characterised by the jet advancing toward the right boundary of the domain with a velocity greater for jet A4 because of the smaller cone angle. Although the shape of the jet is similar, the cone angle modifies the hydrogen concentration, as shown in the figure by . Various experimental investigations confirm this evolution mechanism, and it is characteristic of hollow cone jets Lee, Kim, and Bae (2021); Coratella et al. (2024); Montanaro, Alessandro, Allocca, Luigi, and Meccariello, Giovanni (2024).
@@ V.3 Effects of the nozzle shape
The hollow conical structure of the jet determines a ring-shaped spatial arrangement of the shock waves. The single Mach disc observed with the circular nozzle disappears, but barrel shocks also appear in test cases in A4 and A5; moreover, their size diminishes with the increasing jet angle. The number of these shock cells is larger for case A5. We can conclude that, when performing a hollow cone injection, the increase of the jet angle reduces the Mach number and the hydrogen velocity; therefore, with the same nozzle area, the injection of the same amount of requires a longer time.
Figure 25 shows a volumetric rendering of the vorticity magnitude for the cases A4 and A5 at t = .
To quantitatively assess the mixture quality obtainable with a hollow cone injection, in Figure 26 we plotted the PDF and the spatial average concentration over the domain as a function of time for the cases R2, A4, and A5. To have comparable results, we plotted the PDF at the time instant required to have the same injected mass equal for all the cases.
@@
For all the jets, we can observe a peak of the average YH2 during the first phases of the injection. Almost all the hydrogen in the potential core remains within a few diameters downstream of the exit section, and no mixing occurs despite the intense expansion.
The increased jet angle promotes the mixing with the surrounding air and almost completely removes the hydrogen-rich potential core. The region where the hydrogen mass fraction is close to one is shrunk, for case A5, to a hollow cone around the annular inflow section.
VI Conclusions
We investigated hydrogen under-expanded jets using Large Eddy Simulations of round and annular nozzles, focusing on the turbulent mixing of the hydrogen jets to provide insights and valuable indications to optimise the mixture formation process in low-emission propulsion systems.
We verified and validated the simulations against experimental particle image velocimetry and Schlieren images when available.
The analyses cover different injection pressure ratios and nozzle geometrical configurations. The main outcomes are:
Transient evolution of hydrogen under-expanded jets radically differs from that observed for air. The mixing outside the jet core is enhanced by the baroclinic effect. The structure of the jet core is similar for the two flows in terms of shock structures.
Hydrogen’s physical properties trigger an intense interaction with the surrounding air. The hydrogen jet develops very high vorticity and engulfs a large air volume.
When increasing the pressure ratio, the repeated shock cells disappear, we have a unique Mach disc, and the potential core becomes smaller. The mixture quality slightly improves.
Hollow cone jets have a drastically different morphology when compared to the classical round jets. Ring-shaped shock cells substitute the Mach disc, the jet from highly under-expanded becomes weakly under-expanded, and injecting the same amount of hydrogen requires longer.
The recirculation zone developing downstream of the nozzle for approximately 4D significantly enhances the mixing activity, removes the hydrogen-rich core, and gives a higher quality mixture.
Large jet cone angles (i.e., 135∘) improve the mixture quality from the combustion point of view, quickly reducing the average hydrogen concentration already in the first phase of the injection.
Acknowledgements.
This work has been funded by the European Union - NextGenerationEU, Mission 4, Component 1, under the Italian Ministry of University and Research (MUR) National Innovation Ecosystem grant ECS00000041 - VITALITY - CUP E13C22001060006.Data Availability Statement
The data supporting this study’s findings are available within the article.